OpenStudy (winsteria):

MEDAL AND FAN The area of a rectangle is 41 ft2. The width of the rectangle is 5 ft. What is the length of the rectangle?

OpenStudy (winsteria):

@TrojanPoem

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

8.2 i think

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Yes your correct @ZTS_0216

OpenStudy (anonymous):

Area of a rectangle is length times Width. 41 = 5*x Solve for x.

OpenStudy (trojanpoem):

Area = width * length length = Area/ width = 41/5

OpenStudy (anonymous):

ok thank you @ShadedCookie

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

8.2 is correct because its width * length, so your formula would be 41/5=length

OpenStudy (winsteria):

Thank you

OpenStudy (anonymous):

yw